# Game-Theoretic Coordination For Time-Critical Missions of UAV Systems

### v1

@@ III-B Algorithm
$\begin{split}F_{i}^{ca}\left(s^{k}_{i},\bar{s}^{k}_{-i},\phi\right)=&\sum_{j\neq i}\sum_{\tau=1}^{K}C_{i}\frac{\phi\left(d_{i,j}^{k,\tau},a,b\right)}{\left({d_{i,j}^{k,\tau}}\right)^{2}}.\end{split}$ (27)
@@
$\begin{split}F_{i}^{cm}\left(s^{k}_{i},\bar{s}^{k}_{-i},\phi,\psi\right)=\sum_{j\neq i}\phi\left(d_{i,j}^{k,0},c,d\right)\psi\left(d_{i,j}^{k,0},a,b\right)\\ \times\sum_{\tau=1}^{K}\left(s^{k}_{i\tau}-\bar{s}^{k}_{j\tau}\right)^{2},\end{split}$
where $a<b<c<d$ and
$\psi(z,q_{1},q_{2})=\begin{cases}0,&z<q_{1},\\ \xi(z),&q_{1}\leqslant z\leqslant q_{2},\\ 1,&q_{2}\leqslant z,\end{cases}$ (28)
$q_{1}<q_{2}$ are positive parameters and $0\leqslant\xi(z)\leqslant 1$, such that $\psi\in C^{2}({\mathbb{R}}^{+})$. As a result, time coordination is achieved when the distance between the UAVs is greater than $a$. Furthermore, coordination is disregarded when the distance is less than $a$, and only collision avoidance is active.
Although the cost function (26) is not convex, the optimization problem is still simple enough to solve it efficiently in real-time.
@@ IV Simulations
In this section we demonstrate the effectiveness and applicability of the proposed method in complex, realistic scenarios.
@@ IV-A Simulation Framework
The simulations were performed using RotorPy [12], an open-source Python-based simulator designed for multirotor UAVs. RotorPy provides dynamic modeling of multirotor systems, including aerodynamic effects and nonlinear equations of motion.
The simulation scenarios were designed to explore different aspects of the proposed coordination algorithm. The details of the parameters concerning the MPC used in the simulation scenarios described in this section can be found in Table I. In all scenarios, we use $C_{i}=i+1$ in the collision avoidance term (27). Varying parameters for the scenarios are the communication terms $c$, $d$, and the collision avoidance terms $a$ and $b$, whose values are accordingly presented.
TABLE I: Parameters for MPC used in simulations
Parameters Value
Number of UAVs $N=6$
Prediction horizon $K=10.0$
Time step $h=0.05$ (s)
State constraint $\gamma^{i}_{min}=0.0$
State constraint $\dot{\gamma}^{i}\in\left[0.0,2.0\right]$
Control constraint $\ddot{\gamma}^{i}\in\left[-6.0,6.0\right]$
The algorithm, Alg. 1, employed in these simulations, is designed as follows: the desired state for each UAV is computed, and then passed to the path-following controller that guides the UAV. Subsequently, the path-following error is evaluated, and the initial conditions of the Problem 3 are updated. Then, based on the solution of Problem 3 the UAV transmits $[s^{k}_{i2},\dots,s^{k}_{iK}]$ to the UAVs from the neighbourhood $\mathcal{N}_{ik+1}$. Afterwards, the desired state for the UAV is updated, and the process is repeated iteratively. This process results in the coordination of the multi-agent system.
@@
2Initialize: number of agents $N$, desired trajectories $x_{di}$, state and control constraints $\dot{\gamma}^{i}_{\min}$, $\dot{\gamma}^{i}_{\max}$, $\ddot{\gamma}^{i}_{\max}$, initial conditions $\gamma_{i}^{0}$, $\dot{\gamma_{i}}^{0}$, $i=1,\dots,N$, prediction horizon $K$, total time $T$, time step $h$;
@@ 9            Compute $\alpha_{i}^{k}(x_{i}(kh))$;
10
11            Update initial conditions: $s_{i0}^{k}\leftarrow s^{k-1}_{i1}-\alpha_{i}^{k}$, $l_{i0}^{k}\leftarrow l_{i1}^{k-1}$;
@@
13            $(s_{i}^{k},l_{i}^{k})\leftarrow\textnormal{{SolveProblem3}}(s_{i0}^{k},l_{i0}^{k},\bar{s}_{-i}^{k})$;
@@
15            Transmit $[s^{k}_{i2},\dots,s^{k}_{iK}]$ with UAVs from $\mathcal{N}_{ik+1}$;
@@ IV-A1 Ideal Communication and Ideal Path-Following
We begin with an ideal scenario with perfect communication and no external disturbances, where six UAVs navigate along non-overlapping circular trajectories that share a common center but have different radii. The UAVs aim to minimize the cost function (25), with each starting its mission at a distinct time $\gamma_{1}^{0}=2$, $\gamma_{2}^{0}=1$, $\gamma_{3}^{0}=0$, $\gamma_{4}^{0}=3.5$, $\gamma_{5}^{0}=4$, and $\gamma_{6}^{0}=3$. Note that only the 3rd UAV starts its mission precisely on schedule, as illustrated in Fig. 3(a). The total mission lasts 22 seconds. The results demonstrate rapid synchronization among the UAVs, making $\gamma_{1}=\gamma_{2}=\gamma_{3}=\gamma_{4}=\gamma_{5}=\gamma_{6}$ in around 3.2 seconds, as shown in Fig. 3(b). Moreover, the UAVs successfully adjust their velocities to converge to their desired speed profiles (see Fig. 3(c)). The control input converges to zero by the 7th second (Fig. 4(a)). The simulation results indicate that solutions of Problem 1 are exponentially stable, which validates Theorem 3.
Figure 3: Ideal communication, ideal path following: (a) A top view of the actual trajectories followed by the UAVs under ideal conditions, with red circles marking the starting points and stars indicating the final positions; (b) $\gamma_{i}$ over time; (c) $\dot{\gamma_{i}}$ over time.
@@ IV-A2 Non-Ideal Communication and Ideal Path-Following
Building on the conditions of the ideal communication scenario, this case introduces non-ideal communication among UAVs. The parameters for the communication term are set as follows: $c=3.5$m, $d=7$m. At the beginning of the mission, UAV 5 and UAV 6 are restricted to communicating only with the fourth one. Despite intermittent communication interruptions and subsequent reconnections, the UAVs successfully achieve coordination, with the fourth UAV serving as a communication link between all UAVs.
[…]
@@ IV-A3 Non-Ideal Communication and Non-Ideal Path-Following
In non-ideal path following scenario an additional wind disturbance is introduced along with low communication quality. Wind disturbance affects the motions of UAV 5 and UAV 6 at the beginning of the 22-second long simulation for 10 seconds. The wind force begins with a speed of 5 meters per second and gradually diminishes to zero by the 10-second mark.
The wind-affected trajectories, depicted in Fig. 5(a), show the impact of the path-following term $\alpha_{i}^{k}$ (see (23)) introduced in the algorithm. We set the parameter $\delta=1$ in the path-following error term. This term enables the UAVs to realign with their desired trajectories, ensuring that coordination happens during the mission. However, complete synchronization is observed later compared to the previous scenario, approximately at the 12 second mark, 2 seconds after the wind stops. The control input reduces to zero after the wind effect subsides. In particular, compared to the results of the non-ideal communication and ideal path following simulation results, the drones put more effort in order to adhere to their corresponding desired paths. In the presence of wind disturbance and poor communication quality, the stabilization of the control input is notably delayed, requiring approximately 15 seconds to converge to zero, as illustrated in Fig. 5(c). This represents a significant increase compared to the 10 seconds observed in the ideal path-following scenario, Fig. 4(c).
Figure 5: Non-ideal communication, non-ideal path following: (a) A top view of the trajectories influenced by wind disturbance, with red circles marking the starting points and stars indicating the final positions; (b) $\gamma_{i}$ over time; (c) $\ddot{\gamma_{i}}$ over time.
@@ IV-A4 Collision Avoidance
Based on Remark 2, we demonstrate coordination along with collision avoidance, a scenario involving six UAVs following Lissajous trajectories intersecting at a single central point was implemented (Fig. 6(a)). The trajectories are given by ${\displaystyle x(k)=X\sin(vk+\epsilon)}$ and ${\displaystyle y(k)=Y\sin(wk)}$, with the following parameter values: $X=0.8$, $Y=8$, $v=0.5$, $w=0.25$, $\epsilon=0$. In this setup, no time-delays are implemented; thus $\gamma_{i}^{0}=\gamma_{j}^{0}=0$. The communication terms were set as follows: $c=10$m, $d=20$m, ensuring perfect communication throughout the mission. Collision avoidance parameters were set separately for each of the six UAVs: $a=[2.35,2.5,2.7,2.8,2.9,3.0]$m and $b=[4.7,5.0,5.4,5.6,5.8,6.0]$m. The mission lasted 42 seconds, during which the UAVs completed three passes through their trajectory, successfully avoiding collisions at the intersection point and coordinating afterward. As illustrated in Fig. 6(b), the UAVs avoided collisions between seconds 8 and 13, followed by complete synchronization between seconds 16 and 18. For further clarity, the minimal distance between any two UAVs was plotted over time, Fig. 6(c), which confirms that a minimum separation of at least 0.5 meters was consistently maintained throughout the mission, making sure that no collisions occurred during the simulation.
@@
The method is inherently distributed and relies only on local information, making it naturally scalable. Simulations confirm practical applicability for systems with several dozen UAVs. As the number of agents increases, computation time grows slightly (see Table II), since shared data enters the optimization as constants. However, consensus time decreases, even under partial communication loss, due to increased overall connectivity.
@@
TABLE II: Mean, Max MPC step calculation times and consensus achievement time in communication failure scenario.
@@
The simulation results show the performance of the proposed coordination algorithm in various scenarios. Under ideal conditions, UAVs achieve synchronization rapidly. Introducing non-ideal communication and wind disturbances reveals the ability of the algorithm to handle real-world challenges, ensuring coordination and maintaining safety through collision avoidance mechanisms. For reference, all simulation videos can be found in the supplementary materials.
V Flight Experiments
In this section, we demonstrate the real-time applicability of the method and highlight its communication efficiency, which enables fast information exchange and processing in critical situations. Furthermore, as mentioned before, mission operational requirements can be directly incorporated into the algorithm. In the flight experiments, collision avoidance was integrated as an example.
The flight experiments involve multiple UAVs operating cooperatively in a shared airspace. The results validate the proposed algorithm’s ability to ensure coordination and collision avoidance under realistic conditions of path-following errors and communication failures.
V-1 Non-Ideal Communication and Non-Ideal Path-Following
To validate the effectiveness of the proposed method (see Section III-B), we conducted a flight experiment involving four UAVs exchanging information. Nominal trajectories for the UAVs (Fig. 7(a) dashed circles) are designed to have overlapping regions. The collision avoidance mechanics (see Remark 2) eliminates the requirement for UAV trajectories to maintain spatial separation, as defined in (3). Consequently, UAVs can navigate overlapping paths while ensuring collision-free path-following at all times.
If the distance between UAVs is less than the communication threshold distance ($p_{2}=d$ in (24)), then they exchange information (consensus parameters $s_{i}^{k}$ and desired coordinates $x_{d,i}(s_{i}^{k})$), which enables coordination and collision avoidance via corresponding penalty terms, see (26). Since coordination and collision avoidance represent inherently contradictory objectives—where coordination cannot be maintained if UAVs must simultaneously avoid each other at the coordinates where trajectories intersect—the algorithm dynamically deactivates the coordination penalty term, activates the collision avoidance penalty term when distances between UAVs are smaller than predefined collision avoidance threshold distances ($q_{1}=a$ in (28)). During a collision avoidance maneuver, UAVs are guaranteed to communicate with each other as the collision avoidance threshold distance is always less than the communication threshold distance ($a<d$). If multiple UAVs arrive at a potential collision point with coordinated states and nominal trajectories resemble symmetry, priority is assigned to UAVs with the smallest index (this is achieved by setting $C_{i}=i$ for $i=1,2,\dots,N$ in (27)).
TABLE III: Summary of parameters used for MPC problem (Alg. 1) in flight experiment.
Parameters Value
Number of UAVs $N=4$
Prediction horizon $K=20$
Time step $h=0.2$ (s)
State constraint $\gamma_{min}^{i}=0.0$
State constraint $\dot{\gamma}^{i}\in[0.0,2.0]$
Control constraint $\ddot{\gamma}^{i}\in\left[-15.0,15.0\right]$
Comm. threshold $d=1.0,\,c=0.5$ (m)
Coll. threshold $b=0.39,\,a=0.195$ (m)
V-2 Flight Setup
UAVs follow circular flight paths, the top view of which is shown in Fig. 7(a). The blue, yellow, green, and red curves represent the actual flight trajectories, while the dashed circles represent the nominal trajectories. The UAVs begin their flight on outer circles (black dashed lines) with initial uncoordinated states (black triangles, Table IV) and gradually move to the collision point (blue square) in the clockwise direction to enter the smallest circle (dashed purple circle). After completing a single lap on the smallest circle, all UAVs exit to outer circles from the same collision point (blue square) and continue the flight on outer circles till the final point (purple triangle). Initial states and trajectory parameters are summarized in Table IV. Additionally, $\dot{\gamma}_{i}(0)=1$ for $i=1,2,3,4$, the smallest circle radius, $r_{s}=0.5$ m, the smallest circle center, $\bar{x}_{s}=\left[0.0,0.0,1.0\right]$. Furthermore, the initial velocities, accelerations, angular velocities and yaws are set to zeros, for all UAVs. The circular flight paths were designed to ensure all UAVs have the same angular rates with respect to their circle centers (Table IV), guaranteeing a collision at the collision point and during the smallest circle traversal in the absence of a collision avoidance mechanism. Lastly, we employed the smoothing function (24) within the communication term to attenuate high-frequency effects associated with the second-order derivatives of consensus parameters. This mitigation is essential to prevent potential destabilizing effects on flight performance.
TABLE IV: Initial conditions and trajectory parameters.
Initial Coordination States Initial Positions (m) Outer Circle Radius (m) Outer Circle Center (m)
$\gamma_{1}(0)=0$ $x_{1}(\gamma_{1}(0))=\left[-0.19,-0.7,1.0\right]$ $r_{1}=0.7$ $\bar{x}_{1}=\left[-0.2,0.0,1.0\right]$
$\gamma_{2}(0)=3.33$ $x_{2}(\gamma_{2}(0))=\left[-0.84,-0.77,1.0\right]$ $r_{2}=0.9$ $\bar{x}_{2}=\left[-0.4,0.0,1.0\right]$
$\gamma_{3}(0)=6.66$ $x_{3}(\gamma_{3}(0))=\left[-1.55,-0.55,1.0\right]$ $r_{3}=1.1$ $\bar{x}_{3}=\left[-0.6,0.0,1.0\right]$
$\gamma_{4}(0)=10.0$ $x_{4}(\gamma_{4}(0))=\left[-2.1,0.0,1.0\right]$ $r_{4}=1.3$ $\bar{x}_{4}=\left[-0.8,0.0,1.0\right]$
V-A Experimental Hardware and Software Setup
In this subsection, we review the hardware and software setup, while the final subsection will summarize the experimental results.
The experiment is conducted using four Crazyflie 2.1 mini quadrotors within a 5$\times$7$\times$3m Vicon motion capture arena, utilizing the Crazyswarm software framework [29]. The computational processing for the proposed algorithm is performed on a Ground Station Computer (GSC) equipped with an Intel i7-7700HQ CPU, while the Crazyflies’ integrated PID controllers handle lower-level command execution. The MPC problem, Problem 3, is solved using the CasADi IPOPT numerical solver [4] with parameters listed in Table V-1. Coordinate and orientation data are processed by the Vicon Ground Station Computer (VGSC) and streamed to GSC at 30 Hz via a wired connection. Upon receiving coordinate and orientation information from VGSC, GSC transmits it to the UAVs, along with the state command that UAVs need to follow at 10 Hz, using four Crazyradio 2.0 ISM band radio antennas.
Figure 7: (a) Top view of the flight experiment trajectories. (b) Coordination variable $\gamma_{i}(t)$ for $i=1,2,3,4$, showcasing the temporal evolution of coordination among the UAVs.
V-B Flight Experiments
The results of the flight experiment can be divided into five distinct phases. The first phase spans from $t=0$ to $t=8.6$, during which the UAVs achieve coordination starting from uncoordinated states. This is evident in Fig. 7(b), where the coordination parameters, $\gamma_{i}$, have almost equal values. Fig. 8(b) illustrates how the third and fourth UAVs decelerate ($\ddot{\gamma}_{3},\,\ddot{\gamma}_{4}<0$), while the first and the second UAVs accelerate ($\ddot{\gamma}_{1},\,\ddot{\gamma}_{2}>0$) to catch up in this time interval. Ultimately, the paces of all UAVs converge close to the nominal value ($\dot{\gamma}_{i}=1$) by the end of this phase. The second phase, from $t=8.6$ to $t=19.89$, involves the UAVs performing collision avoidance while transitioning from the outer circles to the smallest circle. This behavior is observable in Fig. 8(b), where the UAVs accelerate at varying rates to avoid collisions. Coordination is temporarily disrupted ($\gamma_{1}\neq\gamma_{2}\neq\gamma_{3}\neq\gamma_{4}$) during this phase, as reflected in Fig. 7(b). The third phase spans $t=19.89$ to $t=37.2$, during which the UAVs continue collision avoidance while gradually reestablishing coordination. By the end of this phase, the pace of the UAVs converges to the nominal values as seen in Fig. 8(a), and the coordination variables stabilize at equal offsets, ensuring collision-free path following. The fourth phase occurs between $t=37.2$ and $t=47.6$, where UAVs avoid collisions while exiting the smallest circle to return to the outer circles. Similar to phase two, the UAVs increase their speeds Fig. 8(a) by accelerating Fig. 8(b) to prevent collisions. In the final fifth phase, spanning $t=47.6$ to $t=58.6$, the UAVs regain coordination on the outer circles. It can be observed in Fig. 7(b), where coordination parameters converge to the same value ($\gamma_{1}=\gamma_{2}=\gamma_{3}=\gamma_{4}$). UAVs complete the mission earlier at $t=58.6$ than the anticipated $t=72.0$ due to maintaining higher-than-nominal pace during collision avoidance maneuvers throughout the flight.
In summary, the outcomes of the flight experiment demonstrate the effectiveness of the proposed algorithm in achieving coordination and collision avoidance among UAVs executing complex time-critical mission in the presence of communication failures. Video footage of this experiment, along with the scenario shown in Fig. 1, is available in the supplementary materials.
TABLE V: Summary of communication between UAVs during the flight experiment.
Time (s) / UAVi,j $z_{1,2}$ $z_{1,3}$ $z_{1,4}$ $z_{2,3}$ $z_{2,4}$ $z_{3,4}$
$2.290<t<t_{1}=$ 2.29
$t_{1}<t<t_{2}=2.89$
$t_{2}<t<t_{3}=14.49$
$t_{3}<t<t_{4}=52.29$
$t_{4}<t<t_{5}=58.69$
Figure 8: (a) $\dot{\gamma}_{i}(t)$ for $i=1,2,3,4$, showcasing the rate of change of the coordination variable over time. (b) $\ddot{\gamma}_{i}(t)$ for $i=1,2,3,4$, showcasing the acceleration of the coordination variable over time.
V-B1 Non-Ideal Communication and Non-Ideal Path-Following Under Wind Disturbance
We also conducted experiment with four UAVs (see Fig. 9) under wind disturbances and communication failures to evaluate the robustness of the proposed algorithm. The fourth UAV experienced the strongest headwind impact. The mean wind speed was 1.34 m/s. The UAV’s ideal speed for coordination in this scenario is 0.37 m/s, while its velocity constraint is 1.4 m/s. The coordination algorithm thus automatically adjusts the speed to enable coordination with minimum deviation.
Figure 9: Flight experiment under wind disturbance with four UAVs.
VI Limitations
In this section, we outline two main limitations of the proposed method. The theoretical analysis of the game-theoretic approach has only been conducted under ideal conditions. Althought the the simulation and experiment results demonstrate the applicability in the case of time-varying networks and the presence of path-following errors; further theoretical developments are necessary to support these results.
In the proposed approach, the desired trajectories are pre- defined and fixed, which restricts the method’s flexibility and applicability. Hence, there is a need to go beyond this assumption and consider trajectory re-planning during mission execution.
VII Conclusion
In this paper, we proposed a novel game-theoretic framework for time-critical cooperative missions of UAV systems operating over a time-varying network. Our approach introduces a distributed time-coordination mechanism that ensures agile, system-wide synchronization in dynamic and uncertain environments. The low dimensionality of the proposed optimization problem allows for effective real-time implementation while incorporating UAV dynamical constraints and mission specifications.
In the simplified scenario, the existence of a Nash equilibrium (system-wide synchronization) and exponential stability of it were proven. Furthermore, to handle more realistic scenarios MPC based algorithm was developed. Extensive simulations demonstrated the method’s effectiveness and reliability in realistic scenarios. Experiments on small UAVs validated the applicability and agility of the method in challenging conditions.
This work represents a reliable and efficient framework for UAV swarm operations, with potential applications in surveillance, disaster management, to name a few.

### v2

@@ III-B1 Collision Avoidance
$\begin{split}G_{i}(s^{k}_{i},\bar{s}^{k}_{-i})=&\sum_{j\neq i}\sum_{\tau=1}^{K}\tfrac{C_{i}\phi(d_{i,j}^{k,\tau},a,b)}{({d_{i,j}^{k,\tau}})^{2}}.\end{split}$ (24)
@@
$F_{i}^{\phi,\psi}(s^{k}_{i},\bar{s}^{k}_{-i})=\sum_{j\neq i}\phi(d_{i,j}^{k,0},c,d)\psi(d_{i,j}^{k,0},a,b)\sum_{\tau=1}^{K}(s^{k}_{i\tau}-\bar{s}^{k}_{j\tau})^{2},$
where $a<b<c<d$ and $\psi(z,q_{1},q_{2})=0$, when $z<q_{1}$, $\psi(z,q_{1},q_{2})=\xi(z)$, when $q_{1}\leqslant z\leqslant q_{2}$ and $\psi(z,q_{1},q_{2})=1,q_{2}\leqslant z$. Here, $q_{1}<q_{2}$ are positive parameters and $0\leqslant\xi(z)\leqslant 1$, such that $\psi\in C^{2}({\mathbb{R}}^{+})$. As a result, time coordination is achieved when the distance between the UAVs is greater than $a$. Furthermore, coordination is disregarded when the distance is less than $a$, and only collision avoidance is active. Although the cost function (23) is not convex, the optimization problem is still simple enough to solve efficiently in real-time.
@@ IV Simulations
In this section, we demonstrate the effectiveness and applicability of the proposed method in complex, realistic scenarios.
@@ IV-A Simulation Framework
The simulations were performed using RotorPy [12], an open-source Python-based simulator designed for multirotor UAVs. In this work, we specifically utilize the Crazyflie multirotor model. RotorPy provides dynamic modeling of multirotor systems, including aerodynamic effects and nonlinear equations of motion.
The simulation scenarios were designed to explore different aspects of the proposed coordination algorithm. The MPC parameters used in all scenarios are as follows: the number of UAVs is fixed at $N=6$ , the prediction horizon is $K=10$, and the time step is $h=0.05\mathrm{~s}$. The state constraints are $\gamma_{\min}^{i}=0$ and $\dot{\gamma}^{i}\in[0,2]$, while the control constraint is $\ddot{\gamma}^{i}\in[-6,6]$. In all simulations, we set $C_{i}=i+1$ in the collision-avoidance term (24). Varying parameters for the scenarios are the communication terms $c$, $d$, and the collision avoidance terms $a$ and $b$, whose values are accordingly presented.
The algorithm, Alg. 1, employed in these simulations, is designed as follows: the desired state for each UAV is computed, and then passed to the path-following controller that guides the UAV. Subsequently, the path-following error is evaluated, and the initial conditions of the Problem 3 are updated. Then, based on the solution of Problem 3 the UAV transmits $s^{k}_{i2},\dots,s^{k}_{iK}]$ to the UAVs from the neighborhood $\mathcal{N}_{ik+1}$. Afterwards, the desired state for the UAV is updated, and the process is repeated iteratively. This process results in the coordination of the multi-agent system.
@@
2Initialize number of agents NN, desired trajectories $x_{di}$, state and control constraints $\dot{\gamma}^{i}_{\min}$, $\dot{\gamma}^{i}_{\max}$, $\ddot{\gamma}^{i}_{\max}$, initial conditions $\gamma_{i}^{0}$, $\dot{\gamma}_{i}^{0}$, $i=1,\dots,N$, prediction horizon $K$, total time $T$, time step $h$;
@@ 8      Compute $\alpha_{i}^{k}(x_{i}(kh))$;
9
10     Update initial conditions;
11    $s_{i0}^{k}\leftarrow s^{k-1}_{i1}-\alpha_{i}^{k}$, $l_{i0}^{k}\leftarrow l_{i1}^{k-1}$;
@@
13    $s_{i}^{k}$, $l_{i}^{k}\leftarrow\textnormal{{SolveProblem3}}(s_{i0}^{k},l_{i0}^{k},\bar{s}_{-i}^{k})$;
@@
15     Transmit $s^{k}_{i2},\dots,s^{k}_{iK}]$ to UAVs from $\mathcal{N}_{i}^{k+1}$;
@@ IV-A1 Ideal Communication and Ideal Path-Following
We consider an ideal scenario with perfect communication and no external disturbances, where all UAVs can exchange information freely. Under these conditions, six UAVs follow non-overlapping, heterogeneous trajectories in three-dimensional space. The UAVs aim to minimize their cost functions, with each starting its mission at a distinct time $\gamma_{1}^{0}=2$, $\gamma_{2}^{0}=1$, $\gamma_{3}^{0}=0$, $\gamma_{4}^{0}=3.5$, $\gamma_{5}^{0}=4$, and $\gamma_{6}^{0}=3$. Note that only the 3rd UAV starts its mission precisely on schedule. The total mission lasts 36 seconds. The results demonstrate rapid synchronization among the UAVs, making $\gamma_{1}=\gamma_{2}=\gamma_{3}=\gamma_{4}=\gamma_{5}=\gamma_{6}$ in around 4.15 seconds, as shown in Fig. 3(b). Moreover, the UAVs successfully adjust their velocities to converge to their desired speed profiles (see Fig. 3(c)). The control input converges to zero by the 6.5th second (Fig. 4(a)). Furthermore, the algorithm’s efficiency is confirmed by a maximum MPC step execution time of 0.0157 seconds, ensuring it does not impose a significant computational burden. The simulation results indicate that solutions of Problem 1 exhibit exponential convergence with respect to the equilibrium trajectory, which validates Theorem 4.
Figure 3: Ideal communication, ideal path following: (a) A top view of the actual trajectories followed by the UAVs under ideal conditions, with circles marking the final positions; (b) $\gamma_{i}$ over time; (c) $\dot{\gamma_{i}}$ over time.
@@ IV-A2 Non-Ideal Communication and Ideal Path-Following
Building on the conditions of the ideal communication scenario, this case introduces non-ideal communication among UAVs, where information exchange is limited. Two UAVs can share their predicted consensus parameters only if the distance between them is below a predefined threshold. The parameters for the communication term are set as follows: c=2.25c=2.25m, d=4.5d=4.5m. At the beggining of the mission, UAV 1 and UAV 2 do not communicate with either UAV 5 or UAV 6. Despite intermittent communication interruptions and subsequent reconnections, the UAVs successfully achieve coordination, with the third and fourth UAVs serving as communication links between all UAVs.
Compared to the ideal communication scenario, system synchronization is achieved in a longer time frame. Specifically, as shown in Fig. 3(b), coordination takes approximately 4.15 seconds in an ideal communication setting. However, under the added challenging communication conditions, the coordination time in this case is almost 11 seconds (see Fig. 4(b)). Along the same lines, the control input converges to zero at around 6.5 seconds in the ideal communication scenario (see Fig. 4(a)), while under non-ideal communication conditions, it takes significantly longer—almost 12.5 seconds—to reach zero, as illustrated in Fig. 4(c). This delay illustrates the impact of poor communication on coordination, as it obstructs the agents’ ability to quickly synchronize and stabilize the control input. Moreover, the presence of the smoothing function $\phi$ in the algorithm does not allow for sharp oscillations of the control input $\ddot{\gamma}$, which, in turn, prevents drastic changes in drone behavior. In this scenario, the maximum time taken for the MPC step calculation is 0.0166.
@@ IV-A3 Non-Ideal Communication and Non-Ideal Path-Following
In a non-ideal path following scenario, an additional wind disturbance is introduced along with partial communication failures. The wind disturbance affects the motions of all UAVs from the start of the 36-second simulation, lasting for a total of 18 seconds (half of the mission time). The disturbance begins with an initial wind speed of 7 m/s that decreases linearly, reaching 0 m/s precisely at the 18-second mark. The wind blows horizontally towards the negative y-axis direction (Fig. 5(a)). This causes the UAVs to significantly shift from their desired trajectories. Additionally, the communication between UAVs is updated every 0.5 seconds, with each pair having a 70% chance of an active communication link.
The wind-affected trajectories, depicted in Fig. 5(a), show the impact of the path-following term $\alpha_{i}^{k}$ (see (21)) introduced in the algorithm. We set the parameter $\delta=1$ in the path-following error term. This term enables the UAVs to realign with their desired trajectories, ensuring that coordination happens during the mission. Synchronization is achieved at the 17.25-second mark, when the wind disturbance dissipates and the UAVs can ideally realign and return to their desired paths. However, communication failures delay the consensus time and affect the control input behavior. In the presence of wind disturbances and partial communication failures, the stabilization of the control inputs is notably delayed, eventually converging to zero at the 25-second mark, as shown in Fig. 5(c). Despite this convergence, residual oscillations are still observed as a clear effect of the communication failures. In such cases, the UAVs exert more effort to stay on their designated paths. The maximum MPC step calculation time is measured to be 0.0192 seconds throughout the simulation.
Figure 5: Non-ideal communication, non-ideal path following: (a) A top view of the trajectories influenced by wind disturbance and partial communication failures, with circles marking the final positions; (b) $\gamma_{i}$ over time; (c) $\ddot{\gamma_{i}}$ over time.
@@ IV-A4 Collision Avoidance
Based on the discussion in Section III-B1, we demonstrate coordination along with collision avoidance, a scenario involving six UAVs following Lissajous trajectories intersecting at a single central point was implemented (Fig. 6(a)). The trajectories are given by $x(k)=X\sin(vk+\epsilon)$} and $y(k)=Y\sin(wk)$}, with the following parameter values: $X=0.8$, $Y=8$, $v=0.5$, $w=0.25$, $\epsilon=0$. In this setup, no time delays are implemented; thus $\gamma_{i}^{0}=\gamma_{j}^{0}=0$. The communication terms were set as follows: c=10c=10m, d=20d=20m, ensuring perfect communication throughout the mission. Collision avoidance parameters were set separately for each of the six UAVs: a=[2.35,2.5,2.7,2.8,2.9,3.0]a=[2.35,2.5,2.7,2.8,2.9,3.0]m and b=[4.7,5.0,5.4,5.6,5.8,6.0]b=[4.7,5.0,5.4,5.6,5.8,6.0]m. The mission lasted 42 seconds, during which the UAVs completed three passes through their trajectory, successfully avoiding collisions at the intersection point and coordinating afterward. As illustrated in Fig. 6(b), the UAVs avoided collisions between seconds 8 and 13, followed by complete synchronization between seconds 16 and 18. For further clarity, the minimal distance between any two UAVs was plotted over time, Fig. 6(c), which confirms that a minimum separation of at least 0.5 meters was consistently maintained throughout the mission, making sure that no collisions occurred during the simulation. During the mission, the maximum execution time for a single MPC step was observed to be 0.028 seconds. This increase in processing time over previous cases is due to the added computational demands of collision avoidance.
@@
The method is inherently distributed and relies only on local information, making it naturally scalable. Simulations confirm practical applicability for systems with several dozen UAVs. To test the scalability of the method, we simulated a scenario in which multiple UAVs fly along non-overlapping circular trajectories that share a common center but have different radii. We assume ideal communication and path-following conditions in this scenario. Under these assumptions, as the number of agents increases, the computation time grows only slightly, as shown in Table II. This limited growth in computation time is due to the fact that information from other agents enters each optimization problem only through fixed coupling terms in the cost function, while the dimensionality of the decision variable remains one.
@@
TABLE II: Mean, Max MPC step calculation times and consensus achievement time in ideal scenario.
@@
The simulation results show the performance of the proposed coordination algorithm in various scenarios. Under ideal conditions, UAVs achieve synchronization rapidly. Introducing non-ideal communication (including partial communication failures) and wind disturbances reveals the ability of the algorithm to handle real-world challenges, ensuring coordination and maintaining safety through collision avoidance mechanisms. Additionally, the method’s inherent scalability, demonstrated through simulations with increasing numbers of UAVs, confirms its practical applicability. For reference, all simulation videos can be found at the following link. The RotorPy simulation code and additional results are available at this link, and simulations at the Crazyswarm platform can be found at this link.
V Conclusion and Future Work
In this paper, we proposed a novel game-theoretic framework for time-critical cooperative missions of UAV systems operating over time-varying networks. The approach introduces a distributed time-coordination mechanism that ensures agile, scalable, system-wide synchronization in dynamic and uncertain environments. The low dimensionality of the proposed optimization problem enables efficient real-time implementation while accommodating UAV dynamical constraints and mission-specific requirements. In the idealized setting, we proved the existence and exponential stability of a Nash equilibrium, ensuring synchronized system behavior. To address more realistic scenarios, we developed an MPC-based algorithm capable of handling communication failures and path-following errors. Extensive simulations demonstrated the effectiveness and robustness of the approach. Although the theoretical analysis has so far been conducted only under ideal conditions, in future work we aim to establish the existence and exponential stability of the Nash equilibrium in more general settings, accounting for time-varying communication networks, potential communication failures, and path-following inaccuracies. These extensions render the cost function nonconvex, while preserving convexity with respect to higher-order derivatives of the virtual time. Consequently, existence of an equilibrium is expected to remain guaranteed, whereas uniqueness may no longer hold. Nevertheless, based on our simulation results, we postulate that all Nash equilibria exhibit exponential convergence properties.
Furthermore, we intend to analyze the convergence properties of the proposed MPC-based algorithm.